Three consecutive odd integers as we know three times the sum of the smaller 2 consecutive odd integers is the same as four times the largest odd integer. Find all three
Answer by macston(5194)

first integer=x; second=x+2; third=x+4
3(x+(x+2))=4(x+4)
6x+6=4x+16 Subtract (4x+6) from each side.
2x=10
x=5 ANSWER 1: The first integer is 5.
x+2=5+2=7 ANSWER 2: The second integer is 7.
x+4=5+4=9 ANSWER 3: The third integer is 9.
CHECK:
3(5+7)=4(9)
3(12)=36
36=36